OpenStudy (anonymous):

Solve the problem, show all the please so i can understand. At many colleges. the number of full time-equivalent students, f is given by f=n/15. where n is the total number of credits for which students enroll in a given Determine the number of full time-equivalent students on a campus in which students registered for a total of 25,350 credits.

OpenStudy (anonymous):

well since its given that total no of credits(n)=25,350 f = n/15 = 25350/15 =1690 so, f=1690 :)